# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ 7. Radial mean bodies
so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely
${\rm R}_{\infty}K=DK=K-K.$
As introduced in Section 2, $\rho_{{\rm R}_{\alpha}K}(u)$ can be expressed as the integral of the chord length $|K\cap(x^{\prime}+\mathbb{R}u)|$. It is also closely related to the covariogram of $K$, that is
$g_{K}(x)=|K\cap(K+x)|.$
In fact
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\rho_{DK}(u)}t^{\alpha-1}g_{K}(tu)dt$
@@
Note that, for $\alpha>0$,
$\rho_{{\rm L}_{\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{K}(x),\operatorname{1}_{K}(x+tu)\}dxdt$
$=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\operatorname{1}_{K}(x)\operatorname{1}_{K}(x+tu)dxdt$
$=\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt$
$=\int_{0}^{\infty}t^{\alpha-1}g_{K}(tu)dt.$
Hence,
${\rm L}_{\alpha}\operatorname{1}_{K}=\Big(\frac{|K|}{\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.1)
for $\alpha>0$. Recall that in Section 2.5,
$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.2)
for $\alpha\in(-1,0)$.
Combining (7.1) and (7.2) with Theorem 2 and Theorem 3, we obtain the following Pólya–Szegö inequality:
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
for $\alpha\in(0,n)$ and
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
for $\alpha>n$. These two formulas can also be regarded as the affine isoperimetric inequalities for chord power integrals.
8. Radial mean bodies for log-concave functions
By (7.1) and (7.2), we define the radial mean bodies for log-concave functions, by
${\rm R}_{\alpha}f=\Big(\frac{-\alpha}{2\|f\|_{1}}\Big)^{1/\alpha}\Pi^{*,-\alpha}f$ (8.1)
for $\alpha\in(-1,0)$ and
${\rm R}_{\alpha}f=\Big(\frac{\alpha}{\|f\|_{1}}\Big)^{1/\alpha}{\rm L}_{\alpha}f$
for $\alpha>0$.
The following proposition describes the relation between ${\rm R}_{\alpha}f$ and the radial $\alpha$-th mean body of the superlevel set $\{f\geq r\}$, for $\alpha\in(-1,0)\cup(0,\infty)$.
Proposition 8.1.
Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. For $\alpha\in(-1,0)\cup(0,\infty)$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),$
where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
@@ 8. Radial mean bodies for log-concave functions
First we consider $\alpha>0$. Recall that
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
where
$\min\{f(x),f(x+tu)\}=\int_{0}^{\|f\|_{\infty}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(x+tu)dr.$
By Fubini’s theorem, we then obtain
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dtdr.$
Note that for a convex body $K$,
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt.$
We have
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}\cdot\frac{|\{f\geq r\}|}{\|f\|_{1}}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
Second, when $\alpha\in(-1,0)$, the definition (8.1) and (2.4) imply that
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$
Since
$|f(x+tu)-f(x)|=\int_{0}^{\infty}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|dr,$
we obtain
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}$ $=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\int_{0}^{\|f\|_{\infty}}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|drdxdt$
$=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dtdr.$
By the calculation in (2.5) and (2.6), we have
$\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}=\frac{-\alpha}{2\{f\geq r\}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dt,$
which implies
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\frac{2|\{f\geq r\}|}{-\alpha}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
This completes the proof. ∎
By the continuity of the $L^{p}$-norm in $p$, we define
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r).$ (8.2)
A more explicit formula for ${\rm R}_{0}f$ will be presented in Section 9.
We recall the following convergence result from [25], which shows that as $\alpha\rightarrow-1$, the radial mean body ${\rm R}_{\alpha}f$ converges to the polar projection body of $f$. The opposite endpoint $\alpha=\infty$ will be considered below.
@@
[25, Theorem 3] For $\alpha\in(-1,0)$ and log-concave $f\in L^{1}(\mathbb{R}^{n})$,
@@
$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u),$
@@
Before proving the lemma below, we recall a result from [18, Theorem 5.5], which states that for every convex body $K$
${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K,$
where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. The complete statement of the theorem will be given in Section 9.
@@
$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha}.$
@@
$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.3)
On the other hand, by the reverse inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$,
$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f>r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha},$
@@
$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{r}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.4)
@@
Applying Jensen’s inequality together with (8.5), we immediately obtain
@@
for $1<\alpha<\beta$. The following theorem provides a functional analogue of (8.6).
@@
If $f$ is not compactly supported, the right inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ is with compact support.
By Proposition 8.1,
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq r\}|\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,r}{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha},$
where $l_{\alpha,r}=((1+\alpha)|\{f\geq r\}|)^{1/\alpha}$ It follows from (8.6) that
$l_{\alpha,r}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)\leq l_{\beta,r}\rho_{{\rm R}_{\beta}\{f\geq r\}}(u),$
@@
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\beta}dr\Big)^{1/\beta}$
@@
Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightharpoonup\infty}(1+\beta)^{1/\beta}=1$, we obtain the right inclusion. Letting $\alpha\rightarrow-1$, by Lemma 8.2, we obtain the left inclusion. ∎
9. Reversed inequalities
In the previous section, we recalled the monotonicity relations (8.5) and (8.6) for radial mean bodies and extended them to their functional analogues. In this section, we turn to the reverse inclusion, which is considerably more delicate.
We first recall the following theorem established by Gardner and Zhang [18]. In particular, for $\alpha=n$,
$c_{n,n}{\rm R}_{n}K\subset n|K|\Pi^{*}K$
which yields Zhang’s inequality [50].

### v2

@@ 5. Radial mean bodies for log-concave functions
so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely,
${\rm R}_{\infty}K={\rm D}K=K-K.$
5.1. Affine isoperimetric inequalities for $\boldsymbol{{\rm R}_{\alpha}f}$
Haddad and Ludwig [24] proved the following affine isoperimetric inequality for ${\rm R}_{\alpha}K$, with full equality characterization.
Theorem 5.1.
[24, Theorem 20] Let $K\subset\mathbb{R}^{n}$ be a convex body. Then
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$
$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n,$
with equality if and only if $K$ is an ellipsoid.
For $\alpha>-1$, the authors [4] defined the radial mean body ${\rm R}_{\alpha}f$ for a non-zero, log-concave function $f\in L^{1}(\mathbb{R}^{n})$ by
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t),~~\alpha\neq 0,$ (5.1)
and
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)d\mu_{f}(t),$ (5.2)
where $d\mu_{f}(t)=\frac{|\{f\geq t\}|}{\|f\|_{1}}dt$ is a probability measure. The following result in [4] shows that ${\rm R}_{\alpha}f$ is proportional to the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$. Moreover, by the continuity of $\rho_{{\rm R}_{\alpha}f}(u)$ with respect to $\alpha$, the authors also gave an explicit formula for $\rho_{{\rm R}_{0}f}$.
Lemma 5.2.
[4, Proposition 6.2, Lemma 6.4] For non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dxdr$ (5.3)
for $\alpha\in(-1,0)$ and
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr$ (5.4)
@@ 5.1. Affine isoperimetric inequalities for $\boldsymbol{{\rm R}_{\alpha}f}$
For $\alpha=0$,
$\log\rho_{{\rm R}_{0}f}(u)=-\gamma+\int_{0}^{\infty}\frac{1}{r}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx-e^{-r}\Big)dr,$
where $\gamma=\Gamma^{\prime}(1)$ is the Euler constant.
The following lemma describes the affine covariance of ${\rm R}_{\alpha}f$, where ${\rm GL}(n)$ denotes the group of general linear transformations on $\mathbb{R}^{n}$.
Lemma 5.3.
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
${\rm R}_{\alpha}(cf)={\rm R}_{\alpha}f~~\text{and}~~{\rm R}_{\alpha}(f\circ\tau_{y})={\rm R}_{\alpha}f,$
where $c>0$ is a constant, $y\in\mathbb{R}^{n}$ and $f\circ\tau_{y}(x)=f(x-y)$. Moreover, for $\phi\in{\rm GL}(n)$,
${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f.$
@@
By (5.3) and (5.4), the radial mean body ${\rm R}_{\alpha}f$ is invariant under scaling and translations. It suffices to show ${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f$.
We treat the case $\alpha>n$. The case $\alpha\in(-1,0)$ can be handled similarly. For $\phi\in{\rm GL}(n)$, note that $\|f\circ\phi^{-1}\|_{1}=|\det\phi|\|f\|_{1}$. Since
$\int_{\mathbb{R}^{n}}\min\{f(\phi^{-1}x),f(\phi^{-1}(x+ru))\}dx=|\det\phi|\int_{\mathbb{R}^{n}}\min\{f(x),f(x+r\phi^{-1}u)\}dx,$
by (5.4), we have
$\rho_{{\rm R}_{\alpha}(f\circ\phi^{-1})}(u)^{\alpha}=\rho_{{\rm R}_{\alpha}f}(\phi^{-1}u)^{\alpha}=\rho_{\phi{\rm R}_{\alpha}f}(u)^{\alpha},$
which concludes the proof. ∎
Recall that the affine chord Sobolev inequalities can be written as
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}},$ $\alpha\in(-1,0),$ (5.5)
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},$ $\alpha\in(0,n)$
and
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},~~\alpha>n.$ (5.6)
By (5.4) and the affine chord Sobolev inequalities, we establish the following affine isoperimetric inequalities for ${\rm R}_{\alpha}f$. We first focus on non-negative $\alpha$.
Note that, when $\alpha=n$ and $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ is log-concave, the chord Sobolev inequalities become the following two trivial inequalities:
$\frac{\|f\|_{1}^{2}}{\|f\|_{\infty}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq\|f\|_{\frac{1}{2}},$ (5.7)
where equality holds if and only if $f=c\operatorname{1}_{K}$ for some convex body $K\subset\mathbb{R}^{n}$ and constant $c>0$. Since
$|{\rm R}_{n}f|=\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy,$
and $|{\rm R}_{n}K|=|K|$ for any convex body $K\subset\mathbb{R}^{n}$, both (5.8) and (5.9) hold for $\alpha=n$ and equality holds if and only if $f=c\operatorname{1}_{K}$.
Theorem 5.4.
Let $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ be log-concave and non-zero. Then
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},\quad\alpha\in(-1,0)\cup(0,n),$ (5.8)
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{1}\|f\|_{\infty}^{-1}}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},\quad\alpha>n.$ (5.9)
In both cases, equality holds if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid.
Proof.
Inequality (5.9) and the equality characterization follow from (5.6) directly. We turn to $\alpha\in(-1,0)\cup(0,n)$. By the affine chord Sobolev inequality (5.5), we obtain
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{n}{n+\alpha}}^{n/\alpha}\|f\|_{1}^{-n/\alpha}}\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|}.$
For brevity, we denote by $\bar{f}(x)=f(x)/\|f\|_{1}$. Then $\bar{f}(x)dx$ is a probability measure on $\mathbb{R}^{n}$. Combining this with the monotonicity of $L^{p}$ norms, for $\alpha\in(-1,n)\setminus\{0\}$, we have
$\|f\|_{\frac{n}{n+\alpha}}^{n/\alpha}\|f\|_{1}^{-n/\alpha}$ $=\Big(\int_{\mathbb{R}^{n}}\bar{f}(x)^{\frac{n}{n+\alpha}}dx\Big)^{\frac{n+\alpha}{\alpha}}$ (5.10)
$=\Big(\int_{\mathbb{R}^{n}}\big(\bar{f}(x)^{-1}\big)^{\frac{\alpha}{n+\alpha}}\bar{f}(x)dx\Big)^{\frac{n+\alpha}{\alpha}}\leq\Big(\int_{\mathbb{R}^{n}}\bar{f}(x)^{\frac{1}{2}}dx\Big)^{2}=\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1},$
where equality holds if and only if $f=c\operatorname{1}_{K}$ for some $c>0$ and convex body $K\subset\mathbb{R}^{n}$. Therefore, we obtain
$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|}.$ (5.11)
If equality holds, then (5.10) implies that $f=c\operatorname{1}_{K}$ where $K\subset\mathbb{R}^{n}$ is a convex body and $c>0$. Hence (5.11) reduces to the geometric inequality (5.1), whose equality cases imply that $K$ is an ellipsoid. ∎
Next we consider the case $\alpha=0$. For non-zero, log-concave $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$, the radial function $\rho_{{\rm R}_{\alpha}f}(u)$ is continuous in $\alpha>-1$, as shown in [4]. Therefore, taking limits on both sides of (5.11) yields the affine isoperimetric inequality for ${\rm R}_{0}f$, but the equality characterization does not follow from this argument. In the following theorem, we use (5.1) together with the definition of ${\rm R}_{0}f$ to establish the inequality with its equality characterization.
Theorem 5.5.
Let $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
$\frac{|{\rm R}_{0}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|},$
where equality holds if and only if $f$ is the characteristic function of an ellipsoid.
Proof.
By (5.2),
$\rho_{{\rm R}_{0}f}(u)^{n}=\exp\Big(\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t)\Big).$
Since $e^{t}$ is a convex function on $\mathbb{R}$ and $\mu_{f}$ is a probability on $\mathbb{R}$, it follows from Jensen’s inequality that
$\rho_{{\rm R}_{0}f}(u)^{n}\leq\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t).$
This, together with (5.1), implies that
$|{\rm R}_{0}f|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm R}_{0}f}(u)^{n}du$ $\leq\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t)du$
$=\int_{0}^{\|f\|_{\infty}}|{\rm R}_{0}\{f\geq t\}|d\mu_{f}(t)$
$\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}\int_{0}^{\|f\|_{\infty}}\frac{|\{f\geq t\}|^{2}}{\|f\|_{1}}dt$
$=\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|f\|_{1}}dxdy.$
By (5.7), we obtain
$\frac{|{\rm R}_{0}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}.$ (5.12)
If equality holds, then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy=\|f\|_{\frac{1}{2}},$
which implies that $f=c\operatorname{1}_{K}$ for some convex body $K\subset\mathbb{R}^{n}$ and $c>0$. Hence (5.12) reduces to the geometric inequality in (5.1). By the equality characterization of (5.1), $K$ is an ellipsoid. ∎
5.2. Monotonicity of $\boldsymbol{{\rm R}_{\alpha}f}$
To consider the monotonicity of radial mean bodies, we recall the following result by Haddad and Ludwig [24].
@@ 5.2. Monotonicity of $\boldsymbol{{\rm R}_{\alpha}f}$
[24, Theorem 3] For $\alpha\in(-1,0)$ and $f\in L^{1}(\mathbb{R}^{n})$ with bounded variation,
@@
We mainly use Lemma 5.6 for log-concave functions $f\in L^{1}(\mathbb{R}^{n})$. By [4, Lemma 5.4], $\mathcal{H}^{n-1}(\partial\{f\geq t\})$ is a integrable function on $\mathbb{R}$. Hence the coarea formula [14, Theorem 5.9], implies that $f$ has bounded variation. The opposite endpoint $\alpha=\infty$ will be considered below.
@@
$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u),$
@@
$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t)\Big)^{1/\alpha}.$
@@
$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u).$ (5.13)
On the other hand, by the reversed inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$ in (1.12),
$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t)\Big)^{1/\alpha},$
@@
$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u).$ (5.14)
@@
This yields the natural definition of ${\rm R}_{\infty}f$ given below.
@@
Applying Jensen’s inequality, together with (5.15), we immediately obtain
@@
for $1<\alpha<\beta<\infty$. The following theorem provides a functional analogue of (5.16).
@@
If $f$ is not compactly supported, the right-most inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ has compact support.
By the definition of ${\rm R}_{\alpha}f$ in (5.1),
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq t\}|\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha}$
$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,t}{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha},$
where $l_{\alpha,t}=((1+\alpha)|\{f\geq t\}|)^{1/\alpha}$. It follows from (5.16) that
$l_{\alpha,t}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)\leq l_{\beta,t}\rho_{{\rm R}_{\beta}\{f\geq t\}}(u),$
@@
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,t}{\rm R}_{\beta}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha}$
$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,t}{\rm R}_{\beta}\{f\geq t\}}(u)^{\beta}dt\Big)^{1/\beta}$
@@
Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightarrow\infty}(1+\beta)^{1/\beta}=1$, we obtain the right-most inclusion. Letting $\alpha\rightarrow-1^{+}$, the left-most inclusion follows from Lemma 5.6. ∎
6. Reverse inclusion for radial mean bodies
In the previous section, we recalled the monotonicity relations (5.15) and (5.16) for radial mean bodies and extended them to their functional analogues. In this section, we turn to the reverse inclusion, which is considerably more delicate. We begin by stating the following theorem established by Gardner and Zhang [18].